# Traveling supersolid stripe patterns in spin-orbit-coupled Bose-Einstein condensates

### v1

@@ IV Properties of moving striped patterns
and the subscript “mix” indicates that expectation values are evaluated with the order parameter (19). When $s_{p}\neq\pm 1$, the polarization vector exhibits a nonzero component in the $xy$ plane, signaling the spontaneous breaking of spin rotational symmetry around the $z$ axis. The azimuthal angle of the polarization is equal to minus the relative phase $\Delta\theta-\Delta\mu\,t/\hbar$ between the two components of the mixture (19). At $t=0$, the direction of the in-plane polarization is set by the arbitrary parameter $-\Delta\theta$, and it subsequently rotates with a frequency determined by the chemical potential difference $\Delta\mu=\mu_{\uparrow}-\mu_{\downarrow}$. This results in a precession of $\langle\bm{\sigma}\rangle_{\mathrm{mix}}$ around the $z$ axis, a phenomenon known in polariton physics as self-induced Larmor precession [82]. This behavior originates from the $g_{ss}$-dependent term in the Gross-Pitaevskii equation (1), which effectively acts as a self-induced Zeeman splitting.
IV.2 Evaluation of the moving stripe order parameter
@@
In the present work, we employ two complementary approaches to determine the order parameter of traveling stripe patterns. The first is a numerical method, which begins by inserting the Ansatz (14) into the total energy (3) and performing the spatial integration. This yields an expression for the energy as a function of the wave vector $k_{1}$ and the expansion coefficients $\tilde{\Psi}_{\bar{m}}$. The energy is then minimized under two constraints enforced via Lagrange multipliers: the chemical potential $\mu$ associated with the normalization condition (16), and the quantity $\Delta\mu/2\equiv\hbar k_{1}v_{x}$ which fixes the value of the momentum (more precisely, the dimensionless ratio $s_{p}=\langle p_{x}\rangle/N\hbar k_{1}$) as required by Eq. (17). Analogously to the zero-$\Omega_{R}$ case discussed in Sec. IV.1, $\Delta\mu$ can here be interpreted as a chemical potential difference between the two minima of the single-particle spectrum and is directly related to the fringe velocity $v_{x}$, while $\mu$ represents the average chemical potential. The stationarity condition of the energy per particle with respect to $k_{1}$, taken at fixed $\mu$, $\Delta\mu$, and $\tilde{\Psi}_{\bar{m}}$, yields the same expression for the optimal wave vector as in the $\bm{v}=0$ case, namely, Eq. (15[58]. To perform the numerical minimization over the components of the $\tilde{\Psi}_{\bar{m}}$’s, the expansion (14) must be truncated to a finite number of harmonics. In our calculations, we retain terms with $-9\leq\bar{m}\leq 9$, which is sufficient to achieve high accuracy in all physical quantities of interest. This choice is validated by direct comparison with the numerical solution of the Gross-Pitaevskii equation (8) in a periodic box, showing excellent agreement.
In Ref. [58], a perturbative approach was developed to analytically compute the equilibrium order parameter in the stripe phase, along with the associated observables. This method has proven highly accurate deeply in the double-minimum regime of the single-particle spectrum, which corresponds to low Raman coupling, i.e., $\hbar\Omega_{R}/4E_{R}\ll 1$. In the present work, we extend this perturbative approach to investigate the case of a moving striped condensate, thereby gaining deeper insight into the underlying physics. This method serves as an analytical alternative to the numerical minimization discussed earlier. Notably, the two approaches yield excellent agreement for small enough values of the Raman coupling (see next section for comparison). We begin by considering the power series expansion of the order parameter,
$\Psi_{0}(\bm{r}^{\prime})=\Psi_{0}^{(0)}(\bm{r}^{\prime})+\sum_{l=1}^{+\infty}\Psi_{0}^{(l)}(\bm{r}^{\prime})\,,$ (21)
along with the expansions for the average chemical potential,
$\mu=G_{dd}+\sum_{l=1}^{+\infty}\mu^{(l)}\,,$ (22)
the chemical potential semi-difference,
$\frac{\Delta\mu}{2}=G_{ss}s_{p}+\sum_{l=1}^{+\infty}\frac{\Delta\mu^{(l)}}{2}\,,$ (23)
and the wave vector,
$k_{1}=k_{R}+\sum_{l=1}^{+\infty}k_{1}^{(l)}\,.$ (24)
Here and in the following, the superscript “$(l)$” denotes the contribution of order $l$ in the small expansion parameter $\hbar\Omega_{R}/4E_{R}$. The zero order terms in each series coincide with the results obtained in the absence of Raman coupling, as discussed in Sec. IV.1. Substituting these expansions into the stationary Gross-Pitaevskii equation (8) yields a set of recurrence relations, which allow for the computation of each order-$l$ correction in terms of lower-order results up to $l-1$. A detailed derivation of this procedure, together with the expressions for the order parameter up to the second order in $\hbar\Omega_{R}/4E_{R}$, is given in Appendix A and Appendix B.
Figure 1: Density profiles of spin-orbit-coupled BEC in the stripe phase for (a) $s_{p}=0$ (stationary stripes), (b) $s_{p}=0.5$, and (c) $s_{p}=0.99$. The numerically obtained profiles of the spin-up (blue solid lines) and spin-down (yellow dashed lines) components are shown. Close to each curve the prediction of the perturbative approach is also plotted (black dotted lines). For definiteness, the fringe offset has been chosen such that a density minimum occurs at $x=0$. The other parameters are $\hbar\Omega_{R}/E_{R}=1.0$, $G_{dd}/E_{R}=1.2$, and $G_{ss}/E_{R}=0.32$.
IV.3 Results for the observables
@@
After evaluating the condensate order parameter using the two methods outlined in the previous section, we now discuss the behavior of several key observables. Throughout this and the following sections, we set the interaction parameters to $G_{dd}/E_{R}=1.2$ and $G_{ss}/E_{R}=0.32$ consistent with the values used in previous works [76, 58]. With this choice of parameters, the first-order transition separating the stripe and plane-wave phases in the ground-state phase diagram occurs at the critical Raman coupling $\hbar\Omega_{\mathrm{cr1}}/E_{R}=2.70$. However, the stripe phase remains metastable up to the spinodal point $\hbar\Omega_{\mathrm{sp}}/E_{R}=2.85$ [58]. In most figures of this work, we compare results for two different values of the Raman coupling. The smaller value, $\hbar\Omega_{R}/E_{R}=1.0$, lies within the range of validity for the perturbative approach, where the predictions of this method match the numerical results excellently. The other value, $\hbar\Omega_{R}/E_{R}=2.6$, is so large that only the numerical method is applicable. It is important to note that the interaction parameters used here differ significantly from those in the original experiment of Ref. [68]. The states of 87Rb employed in that experiment are indeed characterized by a very low degree of miscibility, as evidenced by the small ratio $g_{ss}/g_{dd}\sim 10^{-3}$, resulting in a small critical Raman coupling $\hbar\Omega_{\mathrm{cr1}}/E_{R}=0.19$. This limits the maximum achievable contrast of the stripes and thus restricts the possibility of observing significant supersolidity effects. To address this limitation, potential strategies include using atomic species with tunable interactions, as demonstrated in [22] employing 41K atoms. Alternatively, one could reduce the spatial overlap between the two spin components along a direction parallel to the stripes. This could be achieved by applying a quasi-two-dimensional spin-dependent trapping potential [83, 84, 85] or by implementing spin-orbit coupling between two orbital states within a superlattice potential [86].
@@
We first investigate the effect of stripe motion on the profiles of the total density and spin density. The total density is expressed as
$n(\bm{r}^{\prime})=\bar{n}+\sum_{\bar{m}=1}^{+\infty}\tilde{n}_{\bar{m}}\cos[\bar{m}(2k_{1}x^{\prime}+\Delta\theta)]$ (25)
and the spin density is given by
$s_{z}(\bm{r}^{\prime})=\frac{\langle\sigma_{z}\rangle}{V}+\sum_{\bar{m}=1}^{+\infty}\tilde{s}_{z,\bar{m}}\cos[\bar{m}(2k_{1}x^{\prime}+\Delta\theta)]\,.$ (26)
@@
As with the order parameter in Eq. (14), the total and spin densities are represented as Fourier series. These densities oscillate around their respective average values, $\bar{n}$ and $\langle\sigma_{z}\rangle/V$. The phase $\Delta\theta$ sets the offset of the density fringes. Like in the case of $\Omega_{R}=0$ (see Sec. IV.1), $\Delta\theta$ is randomly selected by the system following the spontaneous breaking of translation invariance. The expansion coefficients $\tilde{n}_{\bar{m}}$ and $\tilde{s}_{z,\bar{m}}$ can be derived directly from those of the order parameter (and the same applies for the spin polarization $\langle\sigma_{z}\rangle$, which is discussed below). At second order in the Raman coupling, only the terms with $\bar{m}=1$ and $\bar{m}=2$ are nonzero. These coefficients are given by
@@
$\begin{split}\frac{\tilde{s}_{z,1}}{\bar{n}}={}&{}2E_{R}\big{[}4E_{R}^{2}G_{ss}+2E_{R}\left(2E_{R}+G_{ss}\right)G_{dd}\\ {}&{}-G_{ss}^{3}s_{p}^{2}\big{]}\frac{s_{p}\sqrt{1-s_{p}^{2}}}{D_{1}}\frac{\hbar\Omega_{R}}{4E_{R}}\,,\end{split}$ (27c)
@@
Here and in the other perturbative formulas presented in this section, the quantities $D_{1}$ and $D_{2}$ appearing in the denominators are even polynomial functions of $s_{p}$. Their full expressions are provided in Appendix B, see Eqs. (59) and (60). The same applies to the numerators $N_{\tilde{n}_{2}}$ and $N_{\tilde{s}_{z,2}}$, given in Eqs. (61) and (62), respectively. It is worth noting that the total density $n$ is an even function of $s_{p}$, while the spin density $s_{z}$ is an odd function. In Fig. 1, we show the combinations $n_{\uparrow,\downarrow}=(n\pm s_{z})/2$, corresponding to the densities of the two spin components, for various values of $s_{p}$. When $s_{p}=0$ [Fig. 1(a)], the spin density vanishes, yielding equal profiles for the two components: $n_{\uparrow}=n_{\downarrow}$. For $s_{p}=0.5$ [Fig. 1(b)], the system becomes both globally and locally spin-polarized, with the two components oscillating around different average values, but remaining in phase. Notably, the minority component (spin-down for $s_{p}>0$, spin-up for $s_{p}<0$) exhibits a larger oscillation amplitude than the majority one. Finally, for $s_{p}=0.99$ [Fig. 1(c)], both components display only weak density modulations around their mean values which closely resemble those of the uniform plane-wave phase.
Figure 2: The fringe contrast [(a1)-(a2)], the fringe wave vector [(b1)-(b2)], and the spin polarization [(c1)-(c2)] as functions of $s_{p}$. In panels [(a1)-(a2)], we show the fringe contrast of the total density (red), as well as the contrast of the majority (blue) and minority (yellow) spin components. Each observable is plotted for two different values of the Raman coupling: $\hbar\Omega_{R}/E_{R}=1.0$ (left panels) and $\hbar\Omega_{R}/E_{R}=2.6$ (right panels). Solid lines correspond to the numerical results. In the left column, close to each of these solid lines, we add a black dotted line showing the predictions of the perturbative approach. The interaction parameters are the same as in Fig. 1: $G_{dd}/E_{R}=1.2$ and $G_{ss}/E_{R}=0.32$.
@@
The contrast of the density profiles of the individual spin components is defined analogously to Eq. (28). As shown in Figs. 2(a1)-(a2), the contrast $\mathcal{C}_{>}$ of the majority component–like that of the total density–is maximal at $s_{p}=0$, decreases monotonically with increasing $|s_{p}|$, and vanishes at $|s_{p}|=1$. In contrast, the behavior of the minority component contrast $\mathcal{C}_{<}$ is nonmonotonic. It initially increases with $|s_{p}|$, reaching the maximum value of $1$ (sooner for larger Raman coupling), and then rapidly drops to zero as $|s_{p}|\to 1$. At first order in the Raman coupling, $\mathcal{C}_{>}$ and $\mathcal{C}_{<}$ can be directly expressed in terms of the coefficients of the expansions (25) and (26). For $s_{p}>0$, where the majority (minority) component corresponds to spin-up (spin-down), one finds $\mathcal{C}_{>}=-(\tilde{n}_{1}+\tilde{s}_{z,1})/[\bar{n}(1+s_{p})]$ and $\mathcal{C}_{<}=-(\tilde{n}_{1}-\tilde{s}_{z,1})/[\bar{n}(1-s_{p})]$. For $s_{p}<0$, the roles of the two components are interchanged. From Figs. 2(a1)-(a2) we observe that the perturbative prediction for $\mathcal{C}_{>}$ matches the numerical results very accurately over the whole range of $|s_{p}|$. However for $\mathcal{C}_{<}$ the perturbative expression is reliable only at small $|s_{p}|$. It eventually diverges unphysically as $|s_{p}|\to 1$. This divergence arises because in the limit $|s_{p}|\to 1$ and for vanishing Raman coupling the denominator in the definition of the contrast [Eq. (28)] tends to zero for the minority component. Consequently, the perturbative approach becomes invalid for calculating $\mathcal{C}_{<}$ at large $|s_{p}|$.
The wave vector $k_{1}$, which determines the periodicity of the stripe phase, also exhibits a dependence on $s_{p}$, albeit much more gradually than the contrast does [see Figs. 2(b1)-(b2)]. This quantity is primarily governed by single-particle physics, with only minor corrections arising from interactions [70, 58]. By substituting the expressions derived in Appendixes into Eq. (15) one obtains the perturbative formula
@@
where the numerator $N_{k_{1}}$ is given in Eq. (63). One can see that $k_{1}$ smoothly interpolates between its value in the stationary stripe phase at $s_{p}=0$ [58] and its value in the plane-wave phase at $|s_{p}|=1$, denoted by $k_{1}^{\mathrm{PW}}$, Eq. (6). Notably, the perturbative result (30) when evaluated at $|s_{p}|=1$ exactly matches the second order expansion of $k_{1}^{\mathrm{PW}}$ in $\hbar\Omega_{R}/4E_{R}$. Finally, we emphasize that $k_{1}$, like the other observables discussed below, should remain invariant under the sign change of the Raman coupling. It is therefore an even function of $\Omega_{R}$.
@@
As already illustrated in Fig. 1, moving stripes exhibit a population imbalance between the spin-up and spin-down components, resulting in a finite spin polarization. In general, the spin polarization $\langle\sigma_{z}\rangle$ is an odd function of $s_{p}$, being positive for $s_{p}>0$ and negative for $s_{p}<0$. Within second order perturbation theory in the Raman coupling, it takes the form
$\langle\sigma_{z}\rangle=Ns_{p}\left[1-\frac{2E_{R}^{2}(2E_{R}+G_{ss})N_{\langle\sigma_{z}\rangle}}{D_{1}^{2}}\left(\frac{\hbar\Omega_{R}}{4E_{R}}\right)^{2}\right]\,,$ (31)
@@
where the coefficient $N_{\langle\sigma_{z}\rangle}$ is defined in Eq. (64). This expression shows that $\langle\sigma_{z}\rangle$ increases linearly with $s_{p}$ for small $|s_{p}|$, in agreement with numerical results plotted in Figs. 2(c1)-(c2). The spin polarization reaches its maximum magnitude at $s_{p}=\pm 1$, where it coincides with the value $Nk_{1}^{\mathrm{PW}}/k_{R}$ characteristic of the plane-wave phase (see Sec. II). It is worth noting that $\langle\sigma_{z}\rangle$, like the stripe wave vector $k_{1}$, decreases as the Raman coupling increases. This behavior reflects the reduced separation between the single-particle energy minima at stronger Raman couplings, which in turn leads to a diminished spin polarization associated with each minimum.
@@
Figure 3: The energy per particle [(a1)-(a2)], the chemical potential [(b1)-(b2)], the chemical potential difference [(c1)-(c2)], and the fringe velocity [(d1)-(d2)] as functions of $s_{p}$. As in Fig. 2, each quantity is shown for two different values of the Raman coupling: $\hbar\Omega_{R}/E_{R}=1.0$ (left panels) and $\hbar\Omega_{R}/E_{R}=2.6$ (right panels). Solid lines represent the results of numerical calculations. In the left panels, close to each of these lines, we add a black dotted line indicating the perturbative predictions. The interaction parameters are the same as in the previous figures: $G_{dd}/E_{R}=1.2$ and $G_{ss}/E_{R}=0.32$.
@@
The perturbative approach also allows for the derivation of analytical expressions for key thermodynamic quantities. The energy per particle, obtained from Eq. (3), reads
$\frac{E}{N}=\frac{G_{dd}}{2}+\frac{G_{ss}}{2}s_{p}^{2}-\frac{E_{R}^{2}N_{E}}{D_{1}}\left(\frac{\hbar\Omega_{R}}{4E_{R}}\right)^{2}\,,$ (32)
while the average chemical potential is given by
$\mu=G_{dd}+\frac{2E_{R}^{2}N_{\mu}}{D_{1}^{2}}\left(\frac{\hbar\Omega_{R}}{4E_{R}}\right)^{2}\,,$ (33)
@@
$\frac{\Delta\mu}{2}=s_{p}\left[G_{ss}-\frac{2E_{R}^{2}(2E_{R}+G_{ss})N_{\Delta\mu}}{D_{1}^{2}}\left(\frac{\hbar\Omega_{R}}{4E_{R}}\right)^{2}\right]\,.$ (34)
The coefficients $N_{E}$, $N_{\mu}$, and $N_{\Delta\mu}$ are given in Eqs. (65), (66), and (67), respectively. These expressions satisfy the thermodynamic relations
$\mu=\left(\frac{\partial E}{\partial N}\right)_{S_{p}}\,,\quad\frac{\Delta\mu}{2}=\left(\frac{\partial E}{\partial S_{p}}\right)_{N}\,,$ (35)
@@
where $S_{p}=Ns_{p}$. Using the identity $v_{x}=\Delta\mu/2\hbar k_{1}$, and combining Eqs. (30) and (34), one obtains the perturbative expression for the fringe velocity,
$v_{x}=\frac{s_{p}}{\hbar k_{R}}\left[G_{ss}-\frac{4E_{R}^{3}N_{v_{x}}}{D_{1}^{2}}\left(\frac{\hbar\Omega_{R}}{4E_{R}}\right)^{2}\right]\,,$ (36)

### v2

@@ IV Properties of moving striped patterns
and the subscript “mix” indicates that expectation values are evaluated with the order parameter (31). When $s_{p}\neq\pm 1$, the polarization vector exhibits a nonzero component in the $xy$ plane, signaling the spontaneous breaking of spin rotational symmetry around the $z$ axis. The azimuthal angle of the polarization is equal to minus the relative phase $\Delta\theta-2\mu_{s}^{(0)}t/\hbar$ between the two components of the mixture (31). At $t=0$, the direction of the in-plane polarization is set by the arbitrary parameter $\Delta\theta$, and it subsequently rotates with a frequency determined by the chemical potential difference $2\mu_{s}^{(0)}=\mu_{\uparrow}-\mu_{\downarrow}$. This results in a precession of $\langle\bm{\sigma}\rangle_{\mathrm{mix}}/V$ around the $z$ axis, a phenomenon known in polariton physics as self-induced Larmor precession [85]. This behavior originates from the $g_{ss}$-dependent term in the Gross-Pitaevskii equation (1), which effectively acts as a self-induced Zeeman splitting.
Figure 2: Density profiles of spin-orbit-coupled BECs in the stripe phase for several values of the population imbalance between dressed spins: (a) $s_{p}=0$ (stationary stripes), (b) $s_{p}=0.5$, and (c) $s_{p}=0.99$. The numerically obtained profiles of the bare spin-up (blue solid lines) and spin-down (yellow dashed lines) components are shown. Close to each curve the prediction of the perturbative approach is also plotted (black dotted lines). For definiteness, the fringe offset has been chosen such that a density minimum occurs at $x^{\prime}=0$. The other parameters are $\hbar\Omega_{R}/E_{R}=1.0$, $G_{dd}/E_{R}=1.2$, and $G_{ss}/E_{R}=0.32$.
IV.2 Results for the observables at finite Raman coupling
@@
After evaluating the condensate order parameter at finite $\Omega_{R}$ using the two methods outlined in Sec. III.3, we now discuss the behavior of several key observables. Throughout this and the following sections, we set the interaction parameters to $G_{dd}/E_{R}=1.2$ and $G_{ss}/E_{R}=0.32$ consistent with the values used in previous works [78, 60]. With this choice of parameters, the first-order transition separating the stripe and plane-wave phases in the ground-state phase diagram occurs at the critical Raman coupling $\hbar\Omega_{\mathrm{cr1}}/E_{R}=2.70$. However, the stripe phase remains metastable up to the spinodal point $\hbar\Omega_{\mathrm{sp}}/E_{R}=2.85$ [60]. In most figures of this work, we compare results for two different values of the Raman coupling. The smaller value, $\hbar\Omega_{R}/E_{R}=1.0$, lies within the range of validity for the perturbative approach, where the predictions of this method match the numerical results excellently. The other value, $\hbar\Omega_{R}/E_{R}=2.6$, is so large that only the numerical method is applicable.
@@
We first investigate the effect of stripe motion on the profiles of the total density and bare spin density. The total density is expressed as
$n(x^{\prime})=\bar{n}+\sum_{\bar{m}=1}^{+\infty}\tilde{n}_{\bar{m}}\cos[\bar{m}(2k_{1}x^{\prime}+\Delta\theta)]$ (33)
and the bare spin density is given by
$s_{z}(x^{\prime})=\bar{s}_{z}+\sum_{\bar{m}=1}^{+\infty}\tilde{s}_{z,\bar{m}}\cos[\bar{m}(2k_{1}x^{\prime}+\Delta\theta)]\,.$ (34)
@@
As with the order parameter in Eq. (19), the total and spin densities are represented as Fourier series. These densities oscillate around their respective average values, $\bar{n}$ and $\bar{s}_{z}$. The phase $\Delta\theta$ sets the offset of the density fringes. Like in the case of $\Omega_{R}=0$ (see Sec. IV.1), $\Delta\theta$ is randomly selected by the system following the spontaneous breaking of translation invariance. The expansion coefficients $\tilde{n}_{\bar{m}}$ and $\tilde{s}_{z,\bar{m}}$ can be derived directly from those of the order parameter (and the same applies for the bare spin polarization density $\bar{s}_{z}$, which is discussed below). At second order in the Raman coupling, only the terms with $\bar{m}=1$ and $\bar{m}=2$ are nonzero. These coefficients are given by
@@
[…]
@@
Here and in the other perturbative formulas presented in this section, the quantities $D_{1}$ and $D_{2}$ appearing in the denominators are even polynomial functions of $s_{p}$. Their full expressions are provided in Appendix B, see Eqs. (73) and (74). The same applies to the numerators $N_{\tilde{n}_{2}}$ and $N_{\tilde{s}_{z,2}}$, given in Eqs. (75) and (76), respectively. It is worth noting that the total density $n$ is an even function of $s_{p}$, while the bare spin density $s_{z}$ is an odd function. In Fig. 2, we show the combinations $n_{\uparrow,\downarrow}=(n\pm s_{z})/2$, corresponding to the densities of the two bare spin components, for various values of $s_{p}$. When $s_{p}=0$ [Fig. 2(a)], the bare spin density vanishes, yielding equal profiles for the two components: $n_{\uparrow}=n_{\downarrow}$. For $s_{p}=0.5$ [Fig. 2(b)], the system becomes both globally and locally spin-polarized, with the two components oscillating around different average values, but remaining in phase. Notably, the minority component (spin-down for $s_{p}>0$, spin-up for $s_{p}<0$) exhibits a larger oscillation amplitude than the majority one. Finally, for $s_{p}=0.99$ [Fig. 2(c)], both components display only weak density modulations around their mean values which closely resemble those of the uniform plane-wave phase.
Figure 3: The fringe contrast [(a1)-(a2)], the mean condensation wave vector [(b1)-(b2)], the half stripe wave vector [(c1)-(c2)], and the bare spin polarization density [(d1)-(d2)] as functions of $s_{p}$. In panels [(a1)-(a2)], we show the fringe contrast of the total density (red), as well as the contrast of the majority (blue) and minority (yellow) bare spin components. Each observable is plotted for two different values of the Raman coupling: $\hbar\Omega_{R}/E_{R}=1.0$ (left panels) and $\hbar\Omega_{R}/E_{R}=2.6$ (right panels). Solid lines correspond to the numerical results. In the left column, close to each of these solid lines, we add a black dotted line showing the predictions of the perturbative approach. The interaction parameters are the same as in Fig. 2: $G_{dd}/E_{R}=1.2$ and $G_{ss}/E_{R}=0.32$.
@@
The contrast of the density profiles of the individual bare spin components is defined analogously to Eq. (36). As shown in Figs. 3(a1)-(a2), the contrast $\mathcal{C}_{>}$ of the majority component, like that of the total density, is maximal at $s_{p}=0$, decreases monotonically with increasing $|s_{p}|$, and vanishes at $|s_{p}|=1$. In contrast, the behavior of the minority component contrast $\mathcal{C}_{<}$ is nonmonotonic. It initially increases with $|s_{p}|$, reaching the maximum value of $1$ (sooner for larger Raman coupling), and then rapidly drops to zero as $|s_{p}|\to 1$. At first order in the Raman coupling, $\mathcal{C}_{>}$ and $\mathcal{C}_{<}$ can be directly expressed in terms of the coefficients of the expansions (33) and (34). For $s_{p}>0$, where the majority (minority) component corresponds to spin-up (spin-down), one finds $\mathcal{C}_{>}=-(\tilde{n}_{1}+\tilde{s}_{z,1})/[\bar{n}(1+s_{p})]$ and $\mathcal{C}_{<}=-(\tilde{n}_{1}-\tilde{s}_{z,1})/[\bar{n}(1-s_{p})]$. For $s_{p}<0$, the roles of the two components are interchanged. From Figs. 3(a1)-(a2) we observe that the perturbative prediction for $\mathcal{C}_{>}$ matches the numerical results very accurately over the whole range of $|s_{p}|$. However for $\mathcal{C}_{<}$ the perturbative expression is reliable only at small $|s_{p}|$. It eventually diverges unphysically as $|s_{p}|\to 1$. This divergence arises because in the limit $|s_{p}|\to 1$ and for vanishing Raman coupling the denominator in the definition of the contrast [Eq. (36)] tends to zero for the minority component. Consequently, the perturbative approach becomes invalid for calculating $\mathcal{C}_{<}$ at large $|s_{p}|$. We point out that, since the density fringes vanish in both bare spin components at $s_{p}=\pm 1$, the system no longer meets the criteria for supersolidity at these points. However, the superfluid fraction remains finite, although smaller than unity [58, 59, 60], indicating that superfluid behavior persists in these limits.
The optimal lengths of the mean condensation wave vector, $k_{c}$, and of the half stripe wave vector, $k_{1}$, also exhibit a dependence on $s_{p}$, albeit much more gradually than the contrast does [see Figs. 3(b1)-(b2) and 3(c1)-(c2)]. These quantities are primarily governed by single-particle physics, which predicts zero $k_{c}$ and $k_{1}$ equal to the wave vector $k_{1}^{\mathrm{SP}}$ characterizing the single-particle energy minima. However, interparticle interactions cause $k_{c}$ to acquire a small but finite value and $k_{1}$ to slightly deviate from $k_{1}^{\mathrm{SP}}$ [72, 60]. Following the procedure illustrated in Appendix A one obtains the perturbative formulas
$k_{c}=-k_{R}s_{p}\frac{4E_{R}^{3}N_{k_{c}}}{D_{1}^{2}}\left(\frac{\hbar\Omega_{R}}{4E_{R}}\right)^{2}$ (38)
and
@@
where the numerators $N_{k_{c}}$ and $N_{k_{1}}$ are given in Eqs. (77) and (78), respectively. One can see that $k_{c}$ is an odd function of $s_{p}$, and takes negative (positive) values for $s_{p}>0$ ($s_{p}<0$). On the other hand, $k_{1}$ is an even function of $s_{p}$ and positive by definition. It is maximum at $s_{p}=0$, where it recovers its value in the stationary stripe phase [60], and minimum at $s_{p}=\pm 1$. We have checked that in the limit $s_{p}\to 1$ the combination $k_{c}+k_{1}$ tends to the condensation wave vector $k_{1}^{\mathrm{PW}}$ characterizing the plane-wave phase, and given by Eq. (10); similarly, one has $k_{c}-k_{1}\to-k_{1}^{\mathrm{PW}}$ as $s_{p}\to-1$. Notably, the sum of the perturbative results (38) and (39) when evaluated at $s_{p}=1$ exactly matches the second order expansion of $k_{1}^{\mathrm{PW}}$ in $\hbar\Omega_{R}/4E_{R}$. Finally, we emphasize that $k_{c}$ and $k_{1}$, like the other observables discussed below, should remain invariant under the sign change of the Raman coupling. They are therefore even functions of $\Omega_{R}$.
@@
As already illustrated in Fig. 2, moving stripes exhibit a population imbalance between the spin-up and spin-down components, resulting in a finite bare spin polarization. In general, the average bare spin polarization density $\bar{s}_{z}$ is an odd function of $s_{p}$, being positive for $s_{p}>0$ and negative for $s_{p}<0$. Within second order perturbation theory in the Raman coupling, it takes the form
$\bar{s}_{z}=\bar{n}s_{p}\left[1-\frac{2E_{R}^{2}(2E_{R}+G_{ss})N_{\bar{s}_{z}}}{D_{1}^{2}}\left(\frac{\hbar\Omega_{R}}{4E_{R}}\right)^{2}\right]\,$, (40)
@@
where the coefficient $N_{\bar{s}_{z}}$ is defined in Eq. (79). This expression shows that $\bar{s}_{z}$ increases linearly with $s_{p}$ for small $|s_{p}|$, in agreement with the numerical results plotted in Figs. 3(d1)-(d2). The average bare spin polarization density reaches its maximum magnitude at $s_{p}=\pm 1$, where it coincides with the value $\bar{n}k_{1}^{\mathrm{PW}}/k_{R}$ characteristic of the plane-wave phase (see Sec. II). It is worth noting that $\bar{s}_{z}$, like the stripe wave vector $k_{1}$, decreases as the Raman coupling increases. This behavior reflects the reduced separation between the single-particle energy minima at stronger Raman couplings, which in turn leads to a diminished bare spin polarization associated with each minimum. We recall that the average momentum density $\bar{p}_{x}$ is proportional to $\bar{s}_{z}$ [see Eq. (26)] and therefore shares the same properties.
@@
Figure 4: The energy per particle [(a1)-(a2)], the mean chemical potential [(b1)-(b2)], the chemical potential semi-difference [(c1)-(c2)], and the fringe velocity [(d1)-(d2)] as functions of $s_{p}$. As in Fig. 3, each quantity is shown for two different values of the Raman coupling: $\hbar\Omega_{R}/E_{R}=1.0$ (left panels) and $\hbar\Omega_{R}/E_{R}=2.6$ (right panels). Solid lines represent the results of numerical calculations. In the left panels, close to each of these lines, we add a black dotted line indicating the perturbative predictions. The interaction parameters are the same as in the previous figures: $G_{dd}/E_{R}=1.2$ and $G_{ss}/E_{R}=0.32$.
@@
The perturbative approach also allows for the derivation of analytical expressions for key thermodynamic quantities. The average energy density, obtained from Eq. (3), reads
$\bar{\varepsilon}=\bar{n}\left[\frac{G_{dd}}{2}+\frac{G_{ss}}{2}s_{p}^{2}-\frac{E_{R}^{2}N_{\bar{\varepsilon}}}{D_{1}}\left(\frac{\hbar\Omega_{R}}{4E_{R}}\right)^{2}\right]\,$, (41)
while the mean chemical potential is given by
$\mu_{d}=G_{dd}+\frac{2E_{R}^{2}N_{\mu_{d}}}{D_{1}^{2}}\left(\frac{\hbar\Omega_{R}}{4E_{R}}\right)^{2}\,$, (42)
@@
$\mu_{s}=s_{p}\left[G_{ss}-\frac{2E_{R}^{2}(2E_{R}+G_{ss})N_{\mu_{s}}}{D_{1}^{2}}\left(\frac{\hbar\Omega_{R}}{4E_{R}}\right)^{2}\right]\,.$ (43)
@@
The coefficients $N_{\bar{\varepsilon}}$, $N_{\mu_{d}}$, and $N_{\mu_{s}}$ are given in Eqs. (80), (81), and (82), respectively. These expressions satisfy the thermodynamic relations (23). Using the identity $v=\mu_{s}/\hbar k_{1}$, and combining Eqs. (39) and (43), one obtains the perturbative expression for the fringe velocity,
$v=\frac{s_{p}}{\hbar k_{R}}\left[G_{ss}-\frac{8E_{R}^{4}N_{v}}{D_{1}^{2}}\left(\frac{\hbar\Omega_{R}}{4E_{R}}\right)^{2}\right]\,$, (44)